# Benjamin Carleton

### v6

@@ 4 Results
[…]
    -$\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{PC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{UW}\leq_{\textnormal{search}}^{p}\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$: Let our input be $(I,S)$. If $S$ is a solution to $I$ for -$\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$, then $I\in B_{\mathcal{E}}$. So output $(\emptyset,C)$. (Note that there is no guarantee on the output if $S$ is not a solution of $I$ for -$\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$ and that is fine since our reduction type does not require us to make any such guarantee. This fact implicitly holds throughout the rest of this proof and so we do not mention it again.)
@@
[…]
@@
[…]
@@
[…]
@@
    $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PV}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}\leq_{\textnormal{search}}^{p}\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PV}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$: The proof of Theorem 3.16 of Carleton et al. [${}^{+}$22] shows how to construct, given $I=(C,V,p)$ and a solution $S$ for $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PV}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$, a solution to $I$ for $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PV}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}$, and we note that the construction can be done in polynomial time.
@@
    $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}\leq_{\textnormal{search}}^{p}\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}$: The proof of Theorem 3.17 of Carleton et al. [${}^{+}$22] shows how to construct, given $I=(C,V,p)$ and a solution $S$ for $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}$, a solution to $I$ for $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}$, and we note that the construction can be done in polynomial time.
    $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}\leq_{\textnormal{search}}^{p}\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}$: The proof of Theorem 3.17 of Carleton et al. [${}^{+}$22] shows how to construct, given $I=(C,V,p)$ and a solution $S$ for $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}$, a solution to $I$ for $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}$, and we note that the construction can be done in polynomial time.
@@
    $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}\leq_{\textnormal{search}}^{p}\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$: The proof of Corollary 3.18 of Carleton et al. [${}^{+}$22] shows how to construct, given $I=(C,V,p)$ and a solution $S$ for $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$, to construct a solution to $I$ for $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$, and we note that the construction can be done in polynomial time.
    $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}\leq_{\textnormal{search}}^{p}\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$: The proof of Corollary 3.18 of Carleton et al. [${}^{+}$22] shows how to construct, given $I=(C,V,p)$ and solution $S$ for $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$, to construct a solution to $I$ for $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$, and we note that the construction can be done in polynomial time.
@@
Proof.    1. This follows directly from Corollaries 4.4 and 4.5 (Parts 2 and 4 of the latter) since approval satisfies Unique-WARP. 2. This follows directly from Corollary 4.5 since approval satisfies Unique-WARP.
We now, for each concrete search-equivalence above about veto and approval, pinpoint the exact search complexity of the search problems.
In the proof of Theorem 4.7 we will draw on the notion of, and known results about, immunity. The notion dates back to the seminal work of Bartholdi, Tovey, and Trick [BTT92] (see also [HHR07], which introduced the destructive cases and fixed a minor flaw in the original formulation for the unique-winner case). We take the definitions’ statements essentially verbatim from Carleton et al. [${}^{+}$22]. In the unique-winner model, we say an election system is immune to a particular type of control if the given type of control can never change a candidate from not uniquely winning to uniquely winning (if the control type is constructive) or change a candidate from uniquely winning to not uniquely winning (if the control type is destructive). In the nonunique-winner model, we say an election system is immune to a particular type of control if the given type of control can never change a nonwinner to a winner (if the control type is constructive) or change a winner to a nonwinner (if the control type is destructive).
@@
Proof.    Consider the case of $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$. Approval is known to be immune to $\mathrm{DC}\hbox{-}\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$ [HHR07]. So on each input $(C,V,p)$, we have: If $p$ is a unique winner of election $(C,V)$ under approval voting, then $p\not\in\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$ (i.e., there exists no candidate partition under which $p$ is not a final-round unique winner in the $\mathrm{PC}\hbox{-}\mathrm{TE}$ two-stage election process under approval voting). Our polynomial-time search-problem algorithm thus is the following: On input $(C,V,p)$, in polynomial time determine whether $p$ is a unique winner under approval of election $(C,V)$. If it is, output $\perp$ (indicating there is no partition that will prevent $p$ from being the unique winner in the $\mathrm{PC}\hbox{-}\mathrm{TE}$ two-stage election process under approval, regarding $(C,V)$). Otherwise output as our solution the partition $(\emptyset,C)$, since this will make the final-round election be $(C,V)$, and we know in this “otherwise” case that $p$ does not uniquely win there.
@@
Since approval is immune to both $\mathrm{CC}\hbox{-}\mathrm{PC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{UW}$ [HHR07] and $\mathrm{CC}\hbox{-}\mathrm{PC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{NUW}$ [${}^{+}$22], the two constructive cases are analogous, of course by asking respectively as to the $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{UW}$ and $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{NUW}$ cases whether $p$ is not a unique winner of $(C,V)$ or not a winner of $(C,V)$, and proceeding in the obvious way, again using $(\emptyset,C)$ as our output partition in those cases that do not output $\perp$.
@@
    Hemaspaandra, Hemaspaandra, and Rothe [HHR07] show that $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PV}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}\in\mathrm{P}$ and their construction detects whether a solution exists and explicitly constructs a solution in polynomial time when a solution exists.
@@
    Hemaspaandra, Hemaspaandra, and Rothe [HHR07] show that $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}\in\mathrm{P}$ and their construction detects whether a solution exists and explicitly constructs a solution in polynomial time when a solution exists.
@@
    Our algorithm, which is a modified version of one of Hemaspaandra, Hemaspaandra, and Rothe [HHR07], proceeds as follows: On input $(C,V,p)$, for each $a\in C$, let $y_{a}$ be the number of votes in $V$ that approve $a$ and let $Y=\max\{y_{a}\mid a\in C\}$. If $y_{p}\neq Y$ and $\|\{a\in C\mid y_{a}=Y\}\|=1$, then output $\perp$. Otherwise, output ($(\{p\},C-\{p\})$. Why does ($(\{p\},C-\{p\})$ work? Note that in this “otherwise” case, we have that either (a) $y_{p}=Y$ or (b) $\|\{a\in C\mid y_{a}=Y\}\|\geq 2$. If (a) holds, then the partition ($(\{p\},C-\{p\})$ makes $p$ a winner in the final round, since $p$ will uniquely win and move forward from its subelection, and no candidate has more approvals than $p$ so even if a candidate moves forward from the other subelection it cannot prevent $p$ from being a final-round winner. If (a) fails and (b) holds, then the partition ($(\{p\},C-\{p\})$ will make $p$ a winner (indeed, a unique winner) in the final round, since in the other subelection there will be at least two candidates that are approved by $Y$ votes, so they tie as winners of that subelection and are eliminated; thus no candidates will move forward from that first-round election.
@@
The following theorem will help us determine the search complexity of the remaining control problems.

### v7

@@ 4.1 Search Equivalences
    Let $\mathcal{E}$ be an election system that satisfies Unique-$\alpha$. The proof of Theorem 3.6 of Carleton et al. [CCH+24] shows that -$\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{PC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{UW}=\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}=B_{\mathcal{E}}=\{(C,V,p)\mid p\in C$ and $p$ is not a unique winner of the $\mathcal{E}$ election (C,V)}(C,V)\}. They also show that for each $I\in B_{\mathcal{E}}$, $(\emptyset,C)$ is a solution to $I$ for both -$\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{PC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{UW}$ and -$\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$.
    -$\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{PC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{UW}\leq_{\textnormal{search}}^{p}\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$: Let our input be $(I,S)$. If $S$ is a solution to $I$ for -$\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$, then $I\in B_{\mathcal{E}}$. So output $(\emptyset,C)$. (Note that there is no guarantee on the output if $S$ is not a solution of $I$ for -$\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$, and that is fine since our reduction type does not require us to make any such guarantee. This fact implicitly holds throughout the rest of this proof and so we do not mention it again.)
@@
    Let $\mathcal{E}$ be an election system that is strongly voiced and that satisfies Property $\alpha$. The proof of Theorem 3.9 of Carleton et al. [CCH+24] shows that -$\mathcal{E}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\mathrm{PC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{UW}=\mathcal{E}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{UW}=A_{\mathcal{E}}=\{(C,V,p)\mid p\in C$ and $p$ is the unique winner of the $\mathcal{E}$ election (C,V)}(C,V)\}. They also show that for each $I\in A_{\mathcal{E}}$, $(\emptyset,C)$ is a solution to $I$ for both -$\mathcal{E}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\mathrm{PC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{UW}$ and -$\mathcal{E}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{UW}$.
@@
    The proof of Theorem 3.12 of Carleton et al. [CCH+24] shows that $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{UW}=\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{UW}=B=\{(C,V,p)\mid p\in C$ and $p$ is not a unique winner of the approval election (C,V)}(C,V)\}. They also show that for each $I\in B$, $(\emptyset,C)$ is a solution to $I$ for both $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{UW}$ and $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{UW}$.
@@
    The proof of Theorem 3.14 of Carleton et al. [CCH+24] shows that $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{NUW}=\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{NUW}=A=\{(C,V,p)\mid p\in C$ and $p$ is a winner of the approval election (C,V)}(C,V)\}. They also show that for each $I\in A$, $(\emptyset,C)$ is a solution to $I$ for both $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{NUW}$ and $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{NUW}$.
@@
    $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PV}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}\leq_{\textnormal{search}}^{p}\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PV}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$: The proof of Theorem 3.16 of Carleton et al. [CCH+24] shows how to construct, given $I=(C,V,p)$ and a solution $S$ for $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PV}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$, a solution to $I$ for $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PV}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}$, and we note that the construction can be done in polynomial time.
@@
    $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}\leq_{\textnormal{search}}^{p}\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}$: The proof of Theorem 3.17 of Carleton et al. [CCH+24] shows how to construct, given $I=(C,V,p)$ and a solution $S$ for $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}$, a solution to $I$ for $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}$, and we note that the construction can be done in polynomial time.
    $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}\leq_{\textnormal{search}}^{p}\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}$: The proof of Theorem 3.17 of Carleton et al. [CCH+24] shows how to construct, given $I=(C,V,p)$ and a solution $S$ for $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}$, a solution to $I$ for $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}$, and we note that the construction can be done in polynomial time.
@@
    $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}\leq_{\textnormal{search}}^{p}\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$: The proof of Corollary 3.18 of Carleton et al. [CCH+24] shows how to construct, given $I=(C,V,p)$ and a solution $S$ for $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$, to construct a solution to $I$ for $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$, and we note that the construction can be done in polynomial time.
    $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}\leq_{\textnormal{search}}^{p}\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$: The proof of Corollary 3.18 of Carleton et al. [CCH+24] shows how to construct, given $I=(C,V,p)$ and solution $S$ for $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$, to construct a solution to $I$ for $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$, and we note that the construction can be done in polynomial time.
@@
Proof. 1. This follows directly from Corollaries 4.4 and 4.5 (Parts 2 and 4 of the latter) since approval satisfies Unique-$\alpha$. 2. This follows directly from Corollary 4.5 since approval is strongly voiced and satisfies Property $\alpha$.
In the next subsection, for each concrete search-equivalence above (i.e., those about plurality, veto, and approval), we pinpoint the exact search complexity of the associated search problems by proving them to be either polynomial-time computable or SAT-equivalent (see Table 1).
4.2 Concrete Search Complexities
In the proof of Theorem 4.7 we will draw on the notion of, and known results about, immunity. The notion dates back to the seminal work of Bartholdi, Tovey, and Trick [BTT92] (see also [HHR07a], which introduced the destructive cases and fixed a minor flaw in the original formulation for the unique-winner case). We take the definitions’ statements essentially verbatim from Carleton et al. [CCH+24]. In the unique-winner model, we say an election system is immune to a particular type of control if the given type of control can never change a candidate from not uniquely winning to uniquely winning (if the control type is constructive) or change a candidate from uniquely winning to not uniquely winning (if the control type is destructive). In the nonunique-winner model, we say an election system is immune to a particular type of control if the given type of control can never change a nonwinner to a winner (if the control type is constructive) or change a winner to a nonwinner (if the control type is destructive).
@@ 4.2 Concrete Search Complexities
Proof.  Consider the case of $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$. Approval is known to be immune to $\mathrm{DC}\hbox{-}\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$ [HHR07a]. So on each input $(C,V,p)$, we have: If $p$ is a unique winner of election $(C,V)$ under approval voting, then $p\not\in\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$ (i.e., there exists no candidate partition under which $p$ is not a final-round unique winner in the $\mathrm{PC}\hbox{-}\mathrm{TE}$ two-stage election process under approval voting). Our polynomial-time search-problem algorithm thus is the following: On input $(C,V,p)$, in polynomial time determine whether $p$ is a unique winner under approval of election $(C,V)$. If it is, output $\perp$ (indicating there is no partition that will prevent $p$ from being the unique winner in the $\mathrm{PC}\hbox{-}\mathrm{TE}$ two-stage election process under approval, regarding $(C,V)$). Otherwise output as our solution the partition $(\emptyset,C)$, since this will make the final-round election be $(C,V)$, and we know in this “otherwise” case that $p$ does not uniquely win there.
@@
Since approval is immune to both $\mathrm{CC}\hbox{-}\mathrm{PC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{UW}$ [HHR07a] and $\mathrm{CC}\hbox{-}\mathrm{PC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{NUW}$ [CCH+24], the two constructive cases are analogous, of course by asking respectively as to the $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{UW}$ and $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{NUW}$ cases whether $p$ is not a unique winner of $(C,V)$ or not a winner of $(C,V)$, and proceeding in the obvious way, again using $(\emptyset,C)$ as our output partition in those cases that do not output $\perp$. ❑
The following theorem is for the most part due to the fact that certain existing [MR18, HHR07a] polynomial-time claims for decision problems are proven via algorithms that are already solving the related search problem that we care about here.
@@
    Hemaspaandra, Hemaspaandra, and Rothe [HHR07a] show that $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PV}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}\in\mathrm{P}$ and their construction detects whether a solution exists and explicitly constructs a solution in polynomial time when a solution exists.
@@
    Hemaspaandra, Hemaspaandra, and Rothe [HHR07a] show that $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}\in\mathrm{P}$ and their construction detects whether a solution exists and explicitly constructs a solution in polynomial time when a solution exists.
@@
    The algorithm of Hemaspaandra, Hemaspaandra, and Rothe [HHR07a] drawn on in the previous part is for the UW model and, not surprisingly, does not solve the NUW case. We now give a modification of that algorithm, and our modified version in fact does, for the NUW case this part is about, construct a solution in polynomial time when a solution exists. Our algorithm proceeds as follows: On input $(C,V,p)$, for each $a\in C$, let $y_{a}$ be the number of votes in $V$ that approve $a$ and let $Y=\max\{y_{a}\mid a\in C\}$. If $y_{p}\neq Y$ and $\|\{a\in C\mid y_{a}=Y\}\|=1$, then output $\perp$. Otherwise, output $(\{p\},C-\{p\})$. Why does $(\{p\},C-\{p\})$ work? Note that in this “otherwise” case, we have that either (a) $y_{p}=Y$ or (b) $\|\{a\in C\mid y_{a}=Y\}\|\geq 2$. If (a) holds, then the partition $(\{p\},C-\{p\})$ makes $p$ a winner in the final round, since $p$ will uniquely win and move forward from its subelection, and no candidate has more approvals than $p$ so even if a candidate moves forward from the other subelection it cannot prevent $p$ from being a final-round winner. If (a) fails and (b) holds, then the partition $(\{p\},C-\{p\})$ will make $p$ a winner (indeed, a unique winner) in the final round, since in the other subelection there will be at least two candidates that are approved by $Y$ votes, so they tie as winners of that subelection and are eliminated; thus no candidates will move forward from that first-round election.
@@
We now turn our attention to proving SAT-equivalence. We start with the following helper theorem.